OpenStudy (anonymous):

How fast will a 2000kg car be moving when it hit the ground if it were dropped from a height 0.2m?

OpenStudy (anonymous):

The important point of this problem is to understand that gravity acts on all objects the same, regardless of their mass. So the fact that the car is 2000 kg is completely irrelevant to answering this. The only information needed is the acceleration of gravity and the initial height. Use the kinematic equation: \[v_{f}^{2}= v_{i}^{2} + 2 a d \rightarrow v_{f}=\sqrt{v_{i}^{2}+2ad}\] Values we have: vi = 0 m/s a = 9.80 m/s^2 d = 0.2 m \[v_{f} = \sqrt{(0)^{2}+(2)(9.80)(0.2)} = \sqrt{3.92}= 1.98\] The car will hit the ground traveling at a speed of 1.98 m/s